OpenStudy (et365):

Another question, 2^(2x+7)=3^(x-13) in terms of logarithmic expression?

OpenStudy (dumbcow):

take log of both sides use property \[\log(x^n) = n \log x\] then simply distribute, add like terms, factor out the x, divide :)

OpenStudy (et365):

Steps after 2xlog(2)+7log(2)=xlog(3)-13log(3)?

OpenStudy (dumbcow):

combine the "x" terms on one side.... constants on other

OpenStudy (dumbcow):

then you will have to factor out the x to get it by itself

OpenStudy (et365):

will -13log(3)-7log(2) become 13log(3)/7log(2) or -13log(3)/7log(2)?

OpenStudy (et365):

I know in this case it might not matter since it is being divided but for future reference.

OpenStudy (dumbcow):

hmm neither ... its just -13log(3) -7log(2)

OpenStudy (et365):

Ahh thank you so much!

OpenStudy (dumbcow):

you may want to change the signs by multiplying everything by a negative

OpenStudy (dumbcow):

yw :)

OpenStudy (et365):

Noted! xD